Funds are classified as winners if they belong in the top 50% by benchmark-adjusted return. The table below shows the percentage of funds that are typically winners in two consecutive years, loser in that period, or the combinations thereto.

What is the probability of a fund being a winner in two consecutive years?

Year 2 winner

Year 2 loser

Year 1 winner

65.5%

34.5%

Year 1 loser

15.5%

84.5%
Answer #1

probability of a fund being a winner in two consecutive years

=P(1st year winner)*P(2nd year winner|1st year winner)=0.5*0.655 =0.3275